# ByteCheckpoint: A Unified Checkpointing System for Large Foundation Model Development

### v2

@@ 5.2 Collective Communication
To enhance communication stability during the planning phase, we have re-implemented the planning procedure using the gRPC framework [14], which eliminates GPU memory usage during planning. However, when scaling training to tens of thousands of GPUs, centralized gathering and scattering operations continued to impose a significant burden on the coordinator, leading to planning communication failures. We further improved the stability of planning by implementing a tree-based hierarchical communication topology. Training workers on a single machine are organized into first-level subtrees, with the worker of local rank 0 designated as the root. For inter-machine communication, we iteratively group multiple machines, designating the worker with the lowest global rank in each group as the root. This procedure continues until all workers are integrated into a hierarchy converging at the global root (i.e., the coordinator). In large-scale 3D parallel training scenarios, a TP-DP-PP communication tree naturally forms, removing extra connections.
Integrity check. A complete checkpoint consists of multiple files stored by different workers. The failure of any single worker can corrupt the entire checkpoint. To prevent such issues, a barrier mechanism is essential to achieve atomic save/load operations among all training workers. Checkpointing modules in training frameworks like Megatron-LM [36] rely on the barrier function in torch.distributed to perform integrity checks. This approach synchronizes training workers to ensure that all checkpoint saving/loading operations are finished. We observed that when scaling training to involving approximately 10,000 GPUs, this behavior results in stalls of about 20 seconds each time. To address this inefficiency, we re-implement the barrier function using the aforementioned method (gRPC with tree-based communication topology) and conduct the integrity checks asynchronously, effectively eliminating the blocking time. We also incorporate upload/download retry mechanisms in ByteCheckpoint’s I/O workers and integrate failure logging, which records the exact stage of failure within the checkpoint saving/loading pipelines in workers who fail to complete checkpointing tasks.
@@ 5.3 Monitoring and Analysis
Figure 12: End-to-end checkpoint saving time heat map from a 3D parallel pre-training task with Megatron-LM on 32 GPUs. The color denotes the duration of selected checkpointing phases on a rank.
Figure 13: Time breakdown of checkpoint saving procedure on rank 0.
@@
Data collection. We have designed a user-friendly metrics system based on Python’s context manager and decorator syntax to flexibly monitor critical procedures. This system automatically captures the duration and I/O size of each operation, along with relevant metadata such as each worker’s rank, the checkpoint file path, and the current training step. All collected metrics are transferred to a remote database via a background message queue.
Visualization. ByteCheckpoint provides users with a comprehensive topological performance overview of all ranks associated with different checkpoint phases, as well as detailed duration breakdowns for any specific rank. Fig. 12 presents an exemplary heat-map visualization of checkpoint saving times within a 3D parallel training topology. The visualization highlights that ranks 0, 4, 8, and 12 experience the longest saving times as their checkpoints including dataloader states. Additional metrics visualizations are also available for detailed analysis, such as blocking time and fine-grained phases like planning and D2H copy. These visualizations enable users to easily pinpoint straggler nodes or phases for both system development and in production environments. For instance, if certain nodes experience network issues, increased upload or download times during the HDFS transfer phase would be readily apparent. Furthermore, detailed timeline breakdowns of checkpointing procedures at each rank are accessible via the heat map topology overview, enabling users to thoroughly evaluate all optimizations in the system. Fig.  13 illustrates the execution details of each checkpoint-saving phase for a specific rank.
Storage-side monitoring. We have also developed a performance and anomaly monitoring system for our storage backends. On the storage client side, we monitor the latency and I/O size of each atomic read/write operation at the I/O chunk level. All collected metrics are periodically transmitted via a message queue to a ClickHouse service, where the data are aggregated and analyzed. Unexpectedly high latency or low bandwidth triggers alerts to engineers for further investigation. On the storage cluster side, our primary focus is on overall performance metrics, including metadata request QPS, cluster-level read/write throughput, and storage capacity utilization. This comprehensive monitoring system plays a crucial role in identifying issues such as slow checkpoint reads and writes, and even complete system unavailability due to storage capacity exhaustion, facilitating the implementation of preventive measures for enhancing system reliability and efficiency.
@@
Model and Framework Source #GPUs Method Blocking Time (s) Saving Time (s) Loading Time (s) Resharding Time (s)
vDiT 4B 32 DCP 16.25 86.82 50.12 74.89
ByteCheckpoint (GPU states) 0.54 (30.09$\times$) 27.47 (3.16$\times$) 11.69 (4.29$\times$) 16.01 (4.68$\times$)
FSDP 128 DCP 61.37 236.34 105.74 91.01
ByteCheckpoint (GPU states) 0.38 (161.50$\times$) 23.74 (9.96$\times$) 12.01 (8.80$\times$) 13.64 (6.67$\times$)
tGPT 70B 2400 MCP 4.73 28.97 69.87 126.30
ByteCheckpoint (GPU states) 0.39 (12.13$\times$) 13.11 (2.21$\times$) 49.48 (1.41$\times$) 62.10 (2.03$\times$)
ByteCheckpoint (full states) 0.50 20.55 72.35 401.21
Megatron-LM 4800 MCP 4.70 76.21 123.80 64.62
ByteCheckpoint (GPU states) 0.36 (13.06$\times$) 8.59 (8.87$\times$) 64.39 (2.00$\times$) 54.31 (1.19$\times$)
ByteCheckpoint (full states) 0.43 25.04 83.56 115.95
6 Experiments
In this section, we detail our deployment and operational experiences with ByteCheckpoint. ByteCheckpoint is built upon DCP [46] (commit hash: 80c07df), a state-of-the-art open-source checkpointing system that supports checkpoint resharding for FSDP [57]. The implementation of ByteCheckpoint consists of about 20,000 lines of Python code, and we are actively working on open-sourcing its components to enhance accessibility and collaboration.
@@ 6.1 Checkpointing Efficiency
Workloads. We adopt two different transformer-based [49] structures, i.e. DiT [31] and GPT-3 [8] to implement two models referred to as vDiT and tGPT, respectively. The vDiT model is fine-tuned with the FSDP framework for video generation, while the tGPT model is trained for text generation with Megatron-LM [36]. Detailed configurations for model and parallelism are provided in Table 3, where the source number of GPUs and parallelism pertain to the saving and loading evaluations, while the target number of GPUs and parallelism correspond to the configurations used for load-time resharding evaluations. All machines in the training clusters are interconnected via InfiniBand, with HDFS selected as the persistent storage solution for all experiments. We integrate the test models into our LFM training trials, conducting training over 500 steps and saving checkpoints every 100 steps. We then resume training under two scenarios: with and without changes in GPU numbers and parallelism configurations, to assess the performance of load-time resharding and standard checkpoint loading. Checkpoint saving performance is evaluated based on average additional overheads, specifically checkpoint stalls (training blocking time) and end-to-end checkpoint saving time (from the save API call to the completion of checkpoint integrity checking). Additionally, we measure the end-to-end checkpoint loading time (blocking time of load API call) for both load-time resharding and standard loading scenarios.
@@
Baselines. In our experiments with FSDP, we use DCP (commit hash: c7338f4) as the baseline , while for Megatron-LM, MCP [40] (commit hash: 3fb5c51) serves as the baseline. We ensure that all training hyperparameters, such as batch size and context length, remain consistent across systems for a fair comparison. Both baselines support asynchronous checkpointing and load-time checkpoint resharding for the model and optimizer, but not for dataloader states. Consequently, we compare I/O performance with these baselines only for GPU states and additionally evaluate performance of ByteCheckpoint for the full training states. Since neither DCP nor MCP supports HDFS, we implement the necessary logic to enable HDFS access for these baselines, using the same configurations (e.g., thread count, file chunk size, etc.) as those employed in ByteCheckpoint.
@@
Saving performance. As shown in Table 4, ByteCheckpoint significantly reduces checkpoint stalls during runtime, with reductions ranging from 13.06$\times$ to 161.50$\times$. This improvement reduces the average checkpoint stall time from minutes or seconds to sub-second durations, outperforming both DCP and MCP. In terms of end-to-end saving time, ByteCheckpoint delivers an average speedup of 6.05$\times$, primarily driven by our fine-grained performance optimizations. Notably, the acceleration provided by ByteCheckpoint scales with the size of the training workload, increasing from 2.21$\times$ for 2400 GPUs to 8.87$\times$ for 4800 GPUs. This can be attributed to the workload balancing mechanism, which becomes more effective with larger DP sizes. Additionally, we observe a particularly significant reduction in blocking time for FSDP workloads. This results from the high communication and synchronization overheads caused by FSDP’s irregular tensor shard processing when using DCP for distributed checkpoint storage. These overheads grow as the training scale increases, making DCP less suitable for large-scale training. In contrast, ByteCheckpoint’s decomposition representation strategy incurs zero communication overhead during metadata generation, leading to improved training efficiency. For training with Megatron-LM, we also report I/O performance for the handling of full training states.
@@
Loading and resharding performance. In experiments involving the loading of distributed checkpoints into unchanged parallelism configurations, ByteCheckpoint achieves performance improvements ranging from 1.41$\times$ to 8.80$\times$ compared to baseline methods. Our approach leverages an asynchronous loading pipeline, which overlaps tensor reading and transferring, significantly reducing blocking time before training begins. When it comes to the resharding efficiency of model and optimizer states, ByteCheckpoint delivers an average acceleration of 3.64$\times$. By eliminating the need for running costly offline resharding jobs, ByteCheckpoint enables flexible and efficient checkpoint transformations across various scenarios. However, we observe that including CPU states, primarily dataloader states, increases the end-to-end load-time resharding time. This is largely due to the time-consuming processing of unique components within the dataloader states, such as the token buffer, which can grow as large as 20GB in text-to-video LFM training. Additionally, since only a subset of workers hold these dataloader states, they often become stragglers during the checkpoint saving and loading (resharding) procedures.
6.2 Microbenchmarks
@@
(a) The normalized loss curve.
(b) The normalized average sample length.
Figure 15: Training resuming with ByteCheckpoint in real production run on 2080 GPUs for several days. Color changes indicate training resuming.
6.3 Correctness Verification
In Fig. 14, we depict normalized training loss curves of tGPT 13B before and after resharding with ByteCheckpoint in various situations. For PP and TP resharding, the loss curve after resharding can smoothly match that in the previous phase. In the case of DP and hybrid resharding, since we also increase the global batch size after resharding, the loss curve displays a consistent but more rapid decline trend. Unit testing confirms that ByteCheckpoint can correctly transform distributed checkpoints across varying parallelism configurations.
We further collect data from real production in Fig. 15, where a 175B language foundation model is trained, and verify the bit-wise alignment ability of ByteCheckpoint when training resumes without changes in parallelism configurations. Since the RNG state is fixed, correct resumptions should yield the same data sampling trajectory. Therefore, we use the normalized sample length curve for evaluation. As highlighted, both the normalized loss and sample length before and after resuming are exactly the same.
Model and Framework #GPUs Parallelism Blocking Time (s) Saving Time (s) Loading Time (s)
Vision Transformer 7B FSDP 1488 ZeRO-2 0.34 20.13 265.73
6.4 Checkpointing in Real LFM Production
Scalabilty. As shown in Table 8, ByteCheckpoint was used to train two transformer-based LFMs for image and text generation, respectively. ByteCheckpoint consistently maintained average checkpoint stalls under 600ms, even at the largest scale with 8,960 GPUs. Additionally, end-to-end checkpointings were completed efficiently, typically within a few seconds.
Dataloader stragglers. By analyzing the performance visualization results from our monitoring tools (Sec. 5.3), we identified prolonged uploading time for dataloader checkpoint files. To support dataloader states resharding, ByteCheckpoint divides the states of a dataloader into multiple parts (e.g., 6 parts for the tGPT experiments in Sec. 6.1) and stores them as individual files. These files can be independently loaded during resharding. However, uploading these small files sequentially leads to low bandwidth utilization, causing performance degradation. In cases like the 7B vision transformer, uploading time of dataloader states can account for 73.16% of total saving time. To fix this issue, we implemented a process pool for concurrent uploads, significantly reducing end-to-end saving time.
@@ 6.4 Checkpointing in Real LFM Production
7 Related Work
Checkpointing frameworks. Some industrial initiatives focus on developing checkpointing systems for deep learning. Prior to DCP [46], PyTorch provided torch.save and torch.load APIs for local checkpoint management but lacked native support for resharding. DCP [46] introduces resharding capabilities for FSDP but does not support parallelism strategies like TP and PP. DeepSpeed-UCP [25] offers a unified format for DeepSpeed checkpoints and provides resharding capabilities through offline scripts. Megatron MCP [40] builds upon the workflow of DCP [46] and extends available file storage formats to include options like Zarr [48]. All these frameworks provide limited support for various parallelisms and training frameworks, and their I/O performance does not scale well in large-scale training.
Checkpointing optimizations. Several works [9, 29] investigated reducing checkpointing costs from different perspectives. Check-N-Run [12], tailored for recommendation models, employs differential checkpointing to store only the modified portions of the model, alongside quantization to reduce checkpoint size. CheckFreq [27] pipelines model states snapshot and saving operations with computation to reduce checkpoint stalls, and designs an online checkpoint frequency tuning algorithm to further reduce cost. Gemini [52] advocates in-memory checkpointing with inter-machine backup for fast failure recovery and interleaves checkpoint communication traffic for state backup with training traffic, enabling frequent checkpointing in each iteration. JIT [15] adopts just-in-time checkpointing for low-cost error recovery. For industrial use cases, it is important to preserve checkpoints to persistent storage due to accessing checkpoints for various purposes like auto-evaluation [18], hyper-parameter tuning, or even restarting training. Different from existing works, ByteCheckpoint champions both optimized I/O performance and flexible automatic online checkpoint resharding, supporting checkpointing for general LFM development.
Checkpoint representation. PyTorch’s torch.save/load functionality relies on pickle for serialization and deserialization. This format lacks crucial tensor shard metadata, such as global shape information, precluding automatic resharding. DCP [46] introduces a disaggregated format that separates metadata from tensor data. The metadata, which includes global shape and offset information, enables DCP’s auto-resharding capabilities. Built upon DCP’s design, BCP incorporates necessary adaptations to support irregular tensor sharding. Array-based storage systems like Zarr [4] and TensorStore [3] allow for saving tensors as individual arrays and support concurrent array reading and writing. MCP [40] supports distributed checkpointing using the Zarr format. For efficient and secure tensor storage, Safetensors [2] is a file foramt for saving tensors safely and loading efficiently. For better compatibility with the Hugging Face open-source ecosystem, ByteCheckpoint incorporates functionality to export checkpoints in the Safetensors format.
Storage systems for LFM. Large-scale checkpointing demands robust storage systems. The LLaMA 3.1 report [11] highlights Tectonic [30], Meta’s general-purpose file system, as the backbone for storing checkpoints during pre-training. One of the primary challenges faced by this system is managing high-frequency checkpoint writes. Similarly, FireFlyer [6], the AI-HPC system developed by DeepSeek [10] for LFM training, incorporates the custom-built 3FS Distributed File System, akin to many parallel file systems such as BeeGFS [17], to handle checkpoint storage.

### v3

@@ 5.2 Collective Communication
To enhance communication stability during planning, we have re-implemented the procedure using the gRPC framework [17], which eliminates GPU memory usage during planning. However, when scaling training to tens of thousands of GPUs, centralized gathering and scattering operations continued to impose a significant burden on the coordinator, leading to communication failures. We further improved its stability by implementing a tree-based hierarchical communication topology. Training workers on a single machine are organized into first-level subtrees, with the worker of local rank 0 designated as the root. For inter-machine communication, we iteratively group multiple machines, designating the worker with the lowest global rank in each group as the root. This procedure continues until all workers are integrated into a hierarchy converging at the global root (i.e., the coordinator). In large-scale 3D parallel training scenarios, a TP-DP-PP communication tree naturally forms, removing extra connections.
@@ 5.3 Monitoring and Analysis
Figure 11: End-to-end checkpoint saving time heat map from a 3D parallel training task with Megatron-LM on 32 GPUs. The color denotes the time of a selected phase on a rank.
Figure 12: Time breakdown of checkpoint saving on rank 0.
@@
Data collection. We have designed a user-friendly metrics system based on Python’s context manager and decorator syntax to monitor critical procedures flexibly. It automatically captures the duration and I/O size of each operation, along with relevant metadata such as each worker’s rank, the file path, and the current step. All collected metrics are transferred to a remote database via a background message queue.
Visualization. ByteCheckpoint provides users with a comprehensive topological performance overview of all ranks associated with different checkpoint phases, as well as detailed duration breakdowns for any specific rank. Fig. 11 presents an exemplary heat-map visualization of checkpoint saving times within a 3D parallel training topology. The visualization highlights that ranks 0, 4, 8, and 12 experience the longest saving times as their checkpoints including dataloader states. Additional metrics visualizations are also available for detailed analysis, such as blocking time and fine-grained phases like planning and D2H copy. These visualizations enable users to easily pinpoint straggler nodes or phases for both system development and in production environments. For instance, if certain nodes experience network issues, increased upload or download times during the HDFS transfer phase would be readily apparent. Furthermore, detailed timeline breakdowns of checkpointing procedures at each rank are accessible via the heat map topology overview, enabling users to thoroughly evaluate all optimizations in the system. Fig.  12 illustrates the execution details of each checkpoint-saving phase for a specific rank.
Storage-side monitoring. On the storage client side, we monitor the latency and I/O size of each atomic read/write operation at the I/O chunk level. All collected metrics are periodically transmitted via a message queue to a ClickHouse service, where the data are aggregated and analyzed. Unexpectedly high latency or low bandwidth triggers alerts to engineers for further investigation. On the storage cluster side, our primary focus is on overall performance metrics, including metadata request QPS, cluster-level read/write throughput, and storage capacity utilization. This comprehensive monitoring system plays a crucial role in identifying issues such as slow checkpoint reads and writes, and even complete system unavailability due to storage capacity exhaustion, facilitating the implementation of preventive measures for enhancing system reliability and efficiency.
@@
Model and Framework Source #GPUs Method $\mathbf{T_{Block}}$ (s) $\mathbf{T_{Save}}$ (s) $\mathbf{T_{Load}}$ (s) $\mathbf{T_{Reshard}}$ (s) ETTR (%)
vDiT 4B 32 DCP 16.25 86.82 50.12 74.89 38.60
ByteCheckpoint (GPU states) 0.54 (30.09$\times$) 27.47 (3.16$\times$) 11.69 (4.29$\times$) 16.01 (4.68$\times$) 46.22 (1.20$\times$)
FSDP 128 DCP 61.37 236.34 105.74 91.01 41.62
ByteCheckpoint (GPU states) 0.38 (161.50$\times$) 23.74 (9.96$\times$) 12.01 (8.80$\times$) 13.64 (6.67$\times$) 48.92 (1.18$\times$)
tGPT 70B 2400 MCP 4.73 28.97 69.87 126.30 34.61
ByteCheckpoint (GPU states) 0.39 (12.13$\times$) 13.11 (2.21$\times$) 49.48 (1.41$\times$) 62.10 (2.03$\times$) 40.18 (1.16$\times$)
ByteCheckpoint (full states) 0.50 20.55 72.35 401.21 28.80
Megatron-LM 4800 MCP 4.70 76.21 123.80 64.62 31.28
ByteCheckpoint (GPU states) 0.36 (13.06$\times$) 8.59 (8.87$\times$) 64.39 (2.00$\times$) 54.31 (1.19$\times$) 40.29 (1.29$\times$)
ByteCheckpoint (full states) 0.43 25.04 83.56 115.95 34.70
6 Evaluation
In this section, we detail our deployment and operational experiences with ByteCheckpoint. ByteCheckpoint is built upon DCP [58] (commit hash: 80c07df), a state-of-the-art open-source checkpointing system that supports checkpoint resharding for FSDP [73]. The implementation of ByteCheckpoint consists of about 20,000 lines of Python code.
@@ 6.1 Checkpointing Efficiency
Workloads. We adopt two different transformer-based [62] structures, i.e. DiT [40] and GPT-3 [9] to implement two models referred to as vDiT and tGPT, respectively. The vDiT model is fine-tuned with the FSDP framework for video generation on a cluster of NVIDIA A100 80GB GPUs, while the tGPT model is trained for text generation with Megatron-LM [47] on NVIDIA H800 80GB GPUs. Detailed configurations for model and parallelism are provided in Table 3, where the source number of GPUs and parallelism pertain to the saving and loading evaluations, while the target number of GPUs and parallelism correspond to the configurations used for load-time resharding evaluations. All machines in the training clusters are interconnected via InfiniBand, with HDFS selected as the persistent storage solution for all experiments. We integrate the test models into our LFM training trials, conducting training over 500 steps and saving checkpoints every 100 steps. We then resume training under two scenarios: with and without changes in GPU numbers and parallelism configurations, to assess the performance of load-time resharding and standard checkpoint loading. Saving performance is evaluated based on average additional overheads, specifically checkpoint stalls (training blocking time) and end-to-end checkpoint saving time (from the save API call to the completion of checkpoint integrity checking). Additionally, we measure the end-to-end checkpoint loading time (the blocking time of the load API call) for both load-time resharding and standard loading scenarios. We use the average ETTR as the metric to evaluate the end-to-end system performance. Following GEMINI [66], we assume a consistent occurrence of failures within each checkpointing interval (100 steps), calculating the average wasted time to derive the achieved ETTR. Calculation details can be found in Appendix C.
@@
Baselines. In experiments with FSDP, we use DCP (commit hash: c7338f4) as the baseline, while for Megatron-LM, MCP [52] (commit hash: 3fb5c51) serves as the baseline. We ensure that all hyper-parameters, such as batch size and context length, remain consistent across systems for a fair comparison. Both baselines support asynchronous checkpointing and load-time checkpoint resharding for the model and optimizer, but not for dataloader states. Consequently, we compare I/O performance with these baselines only for GPU states and additionally evaluate the performance of ByteCheckpoint for the full training states. Since neither DCP nor MCP supports HDFS, we implement the logic to enable HDFS access for them, using the same configurations (e.g., thread count, file chunk size, etc.) as those employed in ByteCheckpoint.
@@
Saving performance. As shown in Table 4, ByteCheckpoint significantly reduces checkpoint stalls during runtime, with reductions ranging from 13.06$\times$ to 161.50$\times$. This improvement reduces the average checkpoint stall time from minutes or seconds to sub-second durations, outperforming both DCP and MCP. In terms of end-to-end saving time, ByteCheckpoint delivers an average speedup of 6.05$\times$, primarily driven by our fine-grained performance optimizations. Notably, the acceleration provided by ByteCheckpoint scales with the size of the training workload, increasing from 2.21$\times$ for 2400 GPUs to 8.87$\times$ for 4800 GPUs. This can be attributed to the workload balancing mechanism, which becomes more effective with larger DP sizes. Additionally, we observe a particularly significant reduction in blocking time for FSDP workloads. This results from the high communication and synchronization overheads caused by FSDP’s irregular tensor shard processing when using DCP for distributed checkpoint storage. These overheads grow as the training scale increases, making DCP less suitable for large-scale training. In contrast, ByteCheckpoint’s decomposition representation strategy incurs zero communication overhead during metadata generation, leading to improved training efficiency. We also provide a detailed overhead breakdown in Appendix D for further investigation. For training with Megatron-LM, we report I/O performance for the handling of full training states.
@@
Loading and resharding performance. In experiments involving the loading of distributed checkpoints into unchanged parallelism configurations, ByteCheckpoint achieves performance improvements ranging from 1.41$\times$ to 8.80$\times$ compared to baseline methods. Our approach leverages an asynchronous loading pipeline, which overlaps tensor reading and transferring, significantly reducing blocking time before training begins. When it comes to the resharding efficiency of model and optimizer states, ByteCheckpoint delivers an average acceleration of 3.64$\times$. By eliminating the need for running costly offline resharding jobs, ByteCheckpoint enables flexible and efficient checkpoint transformations across various scenarios. However, we observe that including CPU states, primarily dataloader states, increases the end-to-end load-time resharding time. This is largely due to the time-consuming processing of unique components within the dataloader states, such as the token buffer, which can grow as large as 20GB in text-to-video LFM training. Additionally, since only a subset of workers hold these dataloader states, they often become stragglers during the checkpoint saving and loading (resharding) procedures. We leave the design of a more efficient dataloader to our future works.
End-to-end performance. ByteCheckpoint outperforms both DCP and MCP in terms of the average end-to-end ETTR, achieving improvements ranging from 1.16$\times$ to 1.29$\times$. Leveraging full-stack optimizations, ByteCheckpoint improves the execution efficiency of each checkpoint-related operation, thereby effectively minimizing training downtime.
@@
Model and Framework #GPUs Parallelism $\mathbf{T_{Block}}$ (s) $\mathbf{T_{Save}}$ (s) $\mathbf{T_{Load}}$ (s)
Vision Transformer 7B FSDP 1488 ZeRO-2 0.34 20.13 265.73
6.2 Microbenchmarks
@@
6.3 Correctness Verification
In Fig. 13, we depict normalized training loss curves of tGPT 13B before and after resharding with ByteCheckpoint in various situations. For PP and TP resharding, the normalized loss curve after resharding can smoothly match that in the previous phase and continue to display a consistent decline trend. Unit testing confirms that ByteCheckpoint can correctly transform distributed checkpoints across varying parallelism. Results of DP and hybrid resharding are in Appendix E.
We further verify the bit-wise alignment ability of ByteCheckpoint when training resumes without changes in parallelism configurations. We show the normalized loss curve from real production in Fig. 14, where a 175B language foundation model is trained. As highlighted, the normalized loss before and after resuming are exactly the same.
6.4 Checkpointing in Real LFM Production
Scalabilty. As shown in Table 8, ByteCheckpoint was used to train two transformer-based LFMs for image and text generation on H800 GPUs, respectively. ByteCheckpoint consistently maintained average checkpoint stalls under 600ms, even at the largest scale with 8,960 GPUs. Additionally, end-to-end checkpointings were completed efficiently, typically within a few seconds.
@@ 6.4 Checkpointing in Real LFM Production
Figure 14: Training resuming with ByteCheckpoint in real production run on 2080 H800 GPUs for several days. Color changes indicate training resuming.
Dataloader stragglers. By analyzing the performance visualization results from our monitoring tools (Sec. 5.3), we identified prolonged uploading time for dataloader checkpoint files. To support dataloader state resharding, ByteCheckpoint divides the states of a dataloader into multiple parts (e.g., 6 parts for the tGPT experiments in Sec. 6.1) and stores them as individual files. These files can be independently loaded during resharding. However, uploading these small files sequentially leads to low bandwidth utilization, causing performance degradation. In cases like the 7B vision transformer, uploading time of dataloader states can account for 73.16% of total saving time. To fix this issue, we implemented a process pool for concurrent uploads, significantly reducing saving time.